OpenStudy (anonymous):

Is x + 5 one of the factors for 4x^2 – 8x – 140 ?

OpenStudy (anonymous):

Yup. You are correct. 4x^2 - 8x - 140 = 4(x^2 - 2x - 35) = 4(x - 7)(x + 5)

OpenStudy (callisto):

You can use the factor theorem to check it. Consider the factor (x+5), f(-5) = 4(-5)^2 – 8(-5) – 140 = 0 Therefore, (x+5) is a factor of 4x^2 – 8x – 140 .